Hi I hope you can help.
My question is 5(2x-1)=4(4-x) I know the answer to this but I cannot figure out how you get to it. The problem I seem to be having is when do two minus signs turn to a plus. I have tried this equation in both ways and can only get the answer if the minus signs change.
Also very confusing was one of the previous questions had two minus signs and would not work unless they remained as minus signs.
So really the question for me is when do you change the signs?
Thankyou very much
Scott Young

Answer by stanbon(75887)

5(2x-1)=4(4-x)
10x-5=16-4x
Add 4x to both sides to get:
14x-5=16
Add 5 to both sides to get:
14x=21
Divide both sides by 14 to get
x=21/14
Reduce 21/14 to get:
x=3/2
---------
Cheers,
Stan H.